OpenStudy (anonymous):

Find the two products below. Compare and contrast, in complete sentences, the similarities and differences of the two. (x + 4)(x − 4) and (x + 4)(x + 4)

jimthompson5910 (jim_thompson5910):

What do you get when you multiply out (x + 4)(x − 4)

OpenStudy (anonymous):

x^2-16?

jimthompson5910 (jim_thompson5910):

and what about (x + 4)(x + 4)

OpenStudy (anonymous):

x^2+8x+16

jimthompson5910 (jim_thompson5910):

good, you are correct on both

jimthompson5910 (jim_thompson5910):

so what do x^2-16 and x^2+8x+16 have in common

OpenStudy (anonymous):

16

jimthompson5910 (jim_thompson5910):

and?

OpenStudy (anonymous):

x^2

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

so both are quadratic expressions

jimthompson5910 (jim_thompson5910):

or polynomials of degree 2

jimthompson5910 (jim_thompson5910):

now onto the differences

OpenStudy (anonymous):

8x

jimthompson5910 (jim_thompson5910):

good, the second expression has that middle term 8x

jimthompson5910 (jim_thompson5910):

so the second expression has 3 terms while the first has 2 terms

jimthompson5910 (jim_thompson5910):

the first is a difference of squares the second is a perfect square

jimthompson5910 (jim_thompson5910):

Another difference: the first is a binomial the second is a trinomial

OpenStudy (anonymous):

thank you!

jimthompson5910 (jim_thompson5910):

oh another similarity is that both are special products...basically they have specific formulas for these cases formulas are: (a-b)(a+b) = a^2 - b^2 ... Difference of squares (a+b)(a+b) = a^2 + 2ab + b^2 ... Perfect Square Trinomial

jimthompson5910 (jim_thompson5910):

you're welcome